[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Jun 26, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Jun 27 02:57:10 UTC 2017


HSU Shift Log for Mon Jun 26 2017 16:28:59 GMT-0400 (EDT)

Shift: H

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None. Satl fixes from SAB are now able to be imported.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: 18Z ECMWF SHIPS intensity guidance did not come in  
despite having been pulled several times between 1505Z-1615Z.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170627/a109217a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list